OpenStudy (anonymous):

Please help me! Simplify by using radical. Rationalize denominators. 2a^-1/2

OpenStudy (anonymous):

why on earth is using radicals called "simplifying"?? please ask your math teacher but \[\large 2a^{-\frac{1}{2}}=\frac{2}{\sqrt{a}}\]

OpenStudy (anonymous):

I do cyber school so I honestly have no idea.

OpenStudy (anonymous):

the exponent \(-\frac{1}{2}\) has a minus sign, which means take the reciprocal a numerator of 1, which is the power, and a denominator oif 2, which means the square root

OpenStudy (anonymous):

that is why \[\large a^{-\frac{1}{2}}=\frac{1}{\sqrt{a}}\]

OpenStudy (anonymous):

This is what it gives me.

OpenStudy (anonymous):



OpenStudy (anonymous):

if you want to get rid of the radical in the denominator you can write it as \[\frac{2\sqrt{a}}{a}\]

OpenStudy (anonymous):

I think that was the answer. Thank you!

OpenStudy (anonymous):

yw

OpenStudy (isaiah.feynman):

@haleyoliviaaa do you understand how he got there?